# ECAMP: Entity-centered Context-aware Medical Vision Language Pre-training

### v3

@@ 5.2 Results on Segmentation and Detection
To assess the effectiveness of local representations learned by ECAMP, extensive experiments on fine-grained downstream tasks including segmentation and detection are conducted in Table. 3. Compared to strong SOTA approaches and boosted by MSCF, ECAMP makes remarkable improvements on all three datasets and ratios. Even in scenarios where MGCA [49] and MedKLIP [54] leverage disease-level labels, and Med-UniC [48] incorporates additional large-scale training data like PadChest [4], our ECAMP consistently outperforms all aforementioned methods. Distinguished Dice score leaps of 4.2% and 1.5% against SOTA methods are achieved on two segmentation datasets SIIM [44] and RSNA [43] using 1% training data, respectively. Notably, ECAMP reaches a Dice score of 67.3%, which outperforms CNN-based methods fine-tuned by 100% training data on SIIM [44]. Similarly, our method attains the best mAP using limited data on the object detection task. Remarkably, our method exhibits a significant lead of 8.3% compared to other SOTA open-source methods based on ViT. These achievements substantiate the effectiveness of ECAMP to optimize informative local features for fine-grained downstream tasks.
@@ 5.3 Ablation Study and Analysis
Advantages of context distillation The first row in Table 4 is trained by basic MIM and MLM. When we use the distilled easy-comprehensible report, we observe a performance improvement of 3.5% on COVIDx [50] and 0.7% on CXR14 [51]. Moreover, when we train ECAMP without the distilled report, notable performance degradation is observed, particularly on COVIDx and RSNA [43], validating the significant role of entity-aware context distillation.
Effectiveness of descriptors masking After constantly masking the critical descriptive words (as shown in Fig 2) preceding the entities, substantial performance enhancements are observed across both coarse-grained and fine-grained tasks, exceeding our initial expectations (as shown in Table 4). Accordingly, it is confirmed that descriptors masking can better leverage the supervised signals in medical reports.
Sensitive analysis of the re-balanced factor In the performance increment between the 3rd and the 4th rows as well as the 7th row and the last row of Table 4, we highlight the effectiveness of our proposed re-balanced factor. The AUC score increases by 0.9% and 0.8% respectively on CXR14, while the ACC shows sequential improvements of 0.7% and 1.5%. It is evident that the re-balanced factor primarily contributes to the performance of classification tasks. We further provide hyper-parameter analysis of $\lambda_{neg}$ corresponding to different imbalance ratios between positive and negative expressions. As shown in Table. 5, setting $\lambda_{neg}$ to $0.05$ according to the imbalance ratio (20:1) leads to best performance.
Performance of context-guided SR As listed in the 4th row and the last row of Table 4, employing context-guided SR provides performance gains from 83.0% to 86.3% on COVIDx. Notably, on SIIM, the Dice score improves from 66.6% to 67.3% and mAP sees a notable increase of 4.9% on RSNA. This underscores the efficacy of context-guided super resolution in aiding pre-trained models to capture prior knowledge of visual texture structure, particularly enhancing performance in fine-grained visual downstream tasks.
@@
Figure 4: Cross-attention maps between diagnosis and vision patterns learned by local aggregated features in MSCF, which successfully localizes the pathology region of various diseases.
Representation analysis As the attention map of the local vision feature and the corresponding distilled reports visualized in Fig. 4, the text modality of ECAMP captures the correct pathological regions of lung collapse (marked by red boxes) through cross-modal attention with the visual modality. Furthermore, there is no particularly disease-sensitive cross-attention for healthy images. These results support the ability of the MSCF to learn accurate and discriminative fine-grained representations, which is crucial for improving the performance of local downstream tasks.
@@ 6 Impact, Limitation, and Conclusion
In this paper, we confront the linguistic challenge of complex medical reports by distilling precise entity-centered knowledge using a large language model. Subsequently, the context-imbalanced issue is addressed by proposing a novel re-balanced factor, then entity-centered descriptors masking is proposed to strengthen the entity-centered context for masked language modeling. Furthermore, a context-guided super-resolution learns to capture the critical pathology vision features with entity-awareness. In addition, multi-scale context fusion helps to optimize informative global and local representations simultaneously. These four readily accessible modules form the core of our simple-yet-effective framework, referred to as ECAMP, whose significant superior performance is validated on various public datasets on downstream tasks, including classification, segmentation, and classification. ECAMP facilitates the development of a robust medical foundation model, ultimately alleviating the workload of experts. However, the absence of explicit alignment between text and vision modalities hinders the application in zero-shot setting. In the future, we plan to incorporate contrastive learning to further empower ECAMP to function as an expert doctor without any manual annotation.
Appendix
Appendix 0.A Statistics of descriptive words
As illustrated in Fig. S1, we present the statistics of our defined descriptive words across 227,835 reports. Specifically, we focus on descriptors with a frequency greater than 10,000 among the total of 31,597 terms. Notably, only "may be" and "there is" imply uncertainty about the existence of the entity, such as a disease or radiology finding. The term "concerning for" can be used to indicate a doctor’s assessment of the presence of a disease based on the interpretation of imaging results. It can also be preceded by "no" to convey the certainty of the absence of a particular disease. With the exception of the descriptive words mentioned above, all other terms listed in Fig. S1 indicate the semantic notion that the entity is absent. Appropriately, the quantity of negative descriptive words is 20×20$\times$20 × greater than that of the positive ones. Thus we propose a rebalanced factor to reweight the loss function to compensate for the dataset bias. Table 5 in the manuscript demonstrates that our re-balanced factor designed according to statistics contributes significantly to the improvements on multiple downstream tasks.
Figure S1: The statistics of descriptive words, here we only show the statistics of descriptors with a frequency greater than 10,000, otherwise, the terms are collectively included in the category labeled as "others".
Appendix 0.B Selected entities
As shown in Table S1, we manually curate a list of 44 entities, including diseases or symptoms, sourced from Medical Subject Headings (MESH [30]). The selection is guided by their importance and the frequency of appearance in radiology reports. Future work involves including more entities, particularly some rare diseases.
Table S1: Our selected entities. We select 44 entities based on their importance and frequency of appearance in radiology reports as comprehensively as possible.
Entities
abnormality abscess aerate aorta
atelectasis bronchiectasis calcification cardiomediastinal
cardiomegaly catheter chf collapse
congestion consolidation contour COPD
deformity dilation distention edema
effusion embolism emphysema engorgement
fibrosis fracture granuloma hernia
hilar hyperinflate hemidiaphragm infiltrate
mass nodule obscure opacity
perihilar pneumonia pneumothorax sarcoidosis
silhouette thickening tuberculosis vasculature
Appendix 0.C Prompt used to distill reports, and examples of distilled reports
Table S2 displays the prompt utilized for report distillation. By inputting the radiology report as the content to ChatGPT [36] with this specific prompt, we obtain a distilled report that includes more detailed semantics regarding the existence and severity of entities. Specifically, using GPT3.5-turbo API [36] setting temperature with 0 to generate 10,000 distilled reports, we train a Vicuna7B [8] to further distill all the reports. Examples of these distilled reports are presented in Table S3.
Table S2: Our prompt used for report distillation is designed to include both an indication and an example. Providing ChatGPT [36] with an example of the task helps enhance its performance in generating distilled reports.
Messages = [ { "role": "user", "content": f""" You are a knowledgeable radiologist. please help me analysis the medical reports and write brief and clear conclusions. Now, I will give you a medical report, as well as the mentioned entities. Please make conclusions according the medical reports, and write brief and clear conclusions as the following format: ’There is [a] [b].’ when you are sure whether the entity exists in the report, or ’There may [a] [b].’, when you are not sure whether the entity exists in the report. [a] represents adjective words describing the severeness of the entity [b]. Please generate conclusions for the entities mentioned above one by one, according to the format and do not generate other words. Please keep only one entity in a sentence, there is no need of using ’and’ or ’or’ to connect two or more words. """ }, { "role": "user", "content": f""" —Example—- Report: as compared to the lung volumes have slightly decreased. signs of mild over inflation and moderate pleural effusion persist. elongation of the descending aorta. Entities: aorta, inflation, effusion Conclusion: There is moderate pleural effusion. There is mild over inflation. There is descending aorta. —Example END—- Given the report: {report} Entities: {entity}. Conclusion: """}]
Table S3: Examples of report distillation are presented. For each report, we showcase how the semantics of our selected entities are distilled by a large language model. The distilled reports provide more specificity regarding the existence of entities.
Original Report
There is no focal consolidation, pleural effusion, vascular congestion or pneumothorax. The cardiac hilar and mediastinal contours within normal limits.
Distilled Report
There is no consolidation. There is no pleural effusion. There is no vascular congestion. There is no pneumothorax. There is normal cardiac hilar. There are mediastinal contours.
Original Report
The lungs are clear of airspace or interstitial opacity. The cardiomediastinal silhouette is unremarkable. No pleural effusion or pneumothorax. No acute or aggressive osseus changes. No evidence of active or latent tb.
Distilled Report
There is no opacity. There is unremarkable cardiomediastinal silhouette. There is no pleural effusion. There is no pneumothorax.
Original Report
Right lower lobe opacity with an associated effusion could represent atelectasis or pneumonia.
Distilled Report
There is opacity. There may be atelectasis or pneumonia in the right lower lobe. There is an associated effusion.
Appendix 0.D Detailed performances on ChestX-ray14
As illustrated in Table S4 and Table S5, we provide a detailed presentation of the results, evaluating the classification performance in both fine-tuning and linear settings. In the fine-tuning setting, our method achieves state-of-the-art results, particularly in atelectasis, consolidation, mass, and pneumonia. However, the performance on edema, emphysema and fibrosis is notably lower. In the linear setting, our method continues to achieve cutting-edge results, especially in consolidation, effusion, and pneumothorax. Similarly, our method faces challenges in edema classification. Future work will prioritize enhancing the classification performance for diseases with suboptimal accuracy.
[…]
Data Ratio
Method
Average
Atelectasis
Cardiomegaly
Consolidation
Edema
Effusion
Emphysema
Fibrosis
Hernia
Infiltration
Mass
Nodule
Pleural Thickening
Pneumonia
Pneumothorax
1% CNN-based
[…]
KAD [60] 78.7 77.0 88.2 82.9 69.2 75.1 69.7 73.5 86.1 72.7 81.3 89.3 74.3 69.2 93.8
ViT-based
REFERS [62] 76.7 77.5 85.6 78.6 84.9 85.4 79.5 72.3 77.1 67.5 76.2 66.5 71.6 69.3 81.7
MRM[63] 79.4 78.8 90.3 80.0 86.5 86.9 82.0 71.9 90.0 67.2 82.3 69.6 72.3 69.6 84.0
ECAMP (Ours) 81.9 80.4 90.3 80.5 88.2 87.4 88.8 77.0 94.1 68.6 85.0 74.0 73.4 70.6 87.8
10% CNN-based
[…]
KAD [60] 80.7 77.6 88.9 83.3 71.8 78.3 71.9 73.7 87.2 75.0 83.3 90.3 80.7 72.3 95.3
ViT-based
REFERS [62] 80.9 80.1 89.8 79.5 87.8 87.5 88.2 77.2 86.1 69.6 82.0 72.8 74.2 72.2 85.6
MRM[63] 84.0 82.3 90.9 81.1 89.0 88.8 92.2 84.8 94.0 70.1 86.6 75.1 78.6 74.3 88.4
ECAMP (Ours) 85.1 83.9 90.9 81.3 89.8 89.1 92.6 86.3 95.6 71.3 87.8 79.0 79.0 76.0 89.3
100% CNN-based
[…]
KAD [60] 82.5 78.5 89.7 84.0 71.3 83.6 77.1 74.0 87.4 75.3 86.0 91.6 82.9 77.8 96.1
ViT-based
REFERS [62] 84.7 83.0 92.3 82.1 90.2 88.7 91.4 83.9 93.3 74.1 85.5 76.7 78.5 77.0 89.1
MRM[63] 85.9 84.2 93.0 82.2 91.0 89.6 94.3 86.7 94.4 71.8 88.2 78.5 81.4 77.3 90.2
ECAMP (Ours) 86.7 85.1 93.1 82.3 91.1 89.7 94.5 87.9 96.0 72.7 88.9 81.4 81.7 78.5 91.0
[…]
Data Ratio
Method
Average
Atelectasis
Cardiomegaly
Consolidation
Edema
Effusion
Emphysema
Fibrosis
Hernia
Infiltration
Mass
Nodule
Pleural Thickening
Pneumonia
Pneumothorax
1% CNN-based
[…]
ViT-based
[…]
ECAMP (Ours) 81.0 80.2 88.4 80.8 88.1 87.8 87.1 76.0 92.2 68.1 82.6 71.4 73.0 71.3 87.6
10% CNN-based
[…]
ViT-based
[…]
ECAMP (Ours) 83.8 82.5 90.6 81.2 89.7 88.9 90.9 84.4 91.2 70.8 85.1 75.6 78.4 75.4 88.8
100% CNN-based
[…]
ViT-based
[…]
ECAMP (Ours) 85.1 83.5 91.5 81.6 90.5 89.2 92.2 86.2 95.6 71.2 85.7 76.6 80.2 77.7 90.1

### v4

@@ 5.2 Results on Segmentation and Detection
To assess the effectiveness of local representations learned by ECAMP, extensive experiments on fine-grained downstream tasks including segmentation and detection are conducted in Table 6. Compared to strong SOTA approaches and boosted by MSCF, ECAMP makes remarkable improvements on all three datasets and ratios. Even in scenarios where MGCA [56] and MedKLIP [61] leverage disease-level labels, and Med-UniC [55] incorporates additional large-scale training data like PadChest [4], our ECAMP consistently outperforms all aforementioned methods. Distinguished Dice score leaps of 4.2% and 1.5% against SOTA methods are achieved on two segmentation datasets SIIM [51] and RSNA [49] using 1% training data, respectively. Notably, ECAMP reaches a Dice score of 67.3%, which outperforms CNN-based methods fine-tuned using all training data on SIIM [51]. Similarly, our method attains the best mAP using limited data on the object detection task. Remarkably, our method exhibits a significant lead of 8.3% compared to other SOTA open-source methods based on ViT. These achievements substantiate the effectiveness of ECAMP to optimize informative local features for fine-grained downstream tasks. Table 11 presents the results of our fine-grained downstream task for retinal cup and disc segmentation on RIGA [20] dataset when pre-trained on FFA images and paired reports. Here, the average Dice coefficient is recorded as the primary metric. Our proposed method, ECAMP, consistently outperforms all previously mentioned approaches on this task, demonstrating its superior performance. Through extensive experiments conducted on FFA images, we have validated that ECAMP effectively leverages its strong capability to utilize contextual information and capture multi-granularity cross-modality relationships between text and images. This enables ECAMP to generalize well across different medical vision tasks, making it a robust and highly effective medical vision-language pre-training method.
[…]
Method RIGA (Dice)
1% 10% 100%
Random init 50.9 65.0 78.3
ImageNet init 63.3 73.8 82.5
ViT-based
MAE [15] 67.0 78.1 86.4
GLoRIA [17] 78.1 82.3 92.4
MRM [73] 83.6 88.3 92.1
ECAMP (Ours) 84.6 88.7 92.6
@@ 5.3 Ablation Study and Analysis
Advantages of context distillation The first row in Table 7 is trained by basic MIM and MLM. When we use the distilled easy-comprehensible report, we observe a performance improvement of 3.5% on COVIDx [57] and 0.7% on ChestX-ray14 [58] while using 1% training data. Moreover, when we train ECAMP without the distilled report, notable performance degradation is observed, particularly on COVIDx and RSNA [49], validating the significant role of entity-aware context distillation.
Effectiveness of descriptors masking After constantly masking the critical descriptive words (as shown in Fig. 3) preceding the entities, substantial performance enhancements are observed across both coarse-grained and fine-grained tasks, exceeding our initial expectations (as shown in Table 7). Accordingly, it is confirmed that descriptors masking can better leverage the supervised signals in medical reports.
Sensitive analysis of the re-balancing factor In the performance increment between the 3rd and the 4th rows as well as the 7th row and the last row of Table 7, we highlight the effectiveness of our proposed re-balancing factor. The AUC score increases by 1.1% and 1.0% respectively on ChestX-ray14 [58], and the ACC shows sequential improvements of 3.5% and 3.2% when utilizing 1% and 10% of training data. It is evident that the re-balancing factor primarily contributes to the performance of classification tasks, and is also effective on fine-grained tasks. We further provide hyper-parameter analysis of $\lambda_{neg}$ corresponding to different imbalance ratios between positive and negative expressions. As shown in Table 9, setting $\lambda_{neg}$ to $0.05$ according to the imbalance ratio (20:1) leads to the best performance. These results strongly demonstrate the effectiveness of re-balanced factor and its capability of encouraging pre-training model to learn precise disease semantics.
Performance of context-guided SR As listed in the 4th row and the last row of Table 7, employing context-guided SR provides performance gains from 83.0% to 86.3% and 92.3% to 95.0% on COVIDx. Notably, on SIIM, the Dice score improves by 0.7% and 0.9%, mAP sees a notable increase of 3.7% and 3.5% on RSNA while using 1% and 10% ratio, which validates its effectiveness in learning informative features by generating detailed pathology vision context. This underscores the efficacy of context-guided super resolution in aiding pre-trained models to capture prior knowledge of visual texture structure, particularly enhancing performance in fine-grained visual downstream tasks.
Fig. 6: Cross-attention maps between diagnosis and vision patterns learned by local aggregated features in MSCF, which successfully localizes the pathology region of various diseases.
@@
Entity selection strategies We conduct systematic ablation studies to evaluate the impact of entity selection strategies. As illustrated in Table 8, we compare three entity selection strategies: (1) Top-20 highest-frequency entities, (2) Top-10% frequency-percentile cutoff, and (3) our proposed method leveraging the Top-44 most frequent entities (as shown in Table 1 ). Compared to our method using 44 selected entities, strategies relying on the Top-20 entities or Top-10% frequency thresholds exhibit diminished performance in modeling disease-related features, particularly under low-data fine-tuning scenarios.
Text input composition As our approach pre-trains vision encoder through the supervision of masked language modeling, text input composition may determine the performance of pre-trained models. For text input composition, we assess four paradigms: (1) original radiology reports alone, (2) ChatGPT-distilled reports alone, (3) hybrid inputs combining original reports with named entity recognition (NER) [19] -based distilled reports, and (4) hybrid inputs combining original reports with ChatGPT-distilled reports. As shown in the last four rows of 8, using ChatGPT-distilled reports alone as input leads to significant performance degradation. While distilled reports capture disease existence and severity, they omit critical visual symptom descriptors, potentially impairing the model’s ability to localize fine-grained anatomical features. Hybrid inputs combining original and distilled reports yield more robust performance, as they balance structured context with raw textual detail.
We conduct further ablation study on context distillation method using the structured template “[entity_1] has [entity_2]” with Radgraph-based NER annotations [19]. Comparing the last two rows of Table 8, NER-based distillation method, though effective for entity presence/absence labeling, is less adept at capturing nuanced clinical relationships in complex reports. This limitation is evidenced by the limited improvement in multi-scale context-aware tasks, especially in fine-grained anatomical location-related tasks (e.g., segmentation and detection). In contrast, ChatGPT-based distillation method preserves nuanced clinical descriptors such as disease severity, anatomical location, and temporal progression (Table 3). Quantitative results demonstrate that our approach more effectively models disease severity and anatomical interactions, strongly supporting our hypothesis that ChatGPT-driven context distillation better preserves implicit clinical knowledge.
5.4 Representation analysis
As the attention map of the local vision feature and the corresponding distilled reports visualized in Fig. 6, the text modality of ECAMP captures the correct pathological regions of lung collapse (marked by red boxes) through cross-modal attention with the visual modality. Furthermore, there is no particularly disease-sensitive cross-attention for healthy images. These results support the ability of the MSCF to learn accurate and discriminative fine-grained representations, which is crucial for improving the performance of local downstream tasks.
@@ 6 Impact, Limitation, and Conclusion
In this paper, we propose a novel pre-training method which extracts representative and multi-scale features from complex and imbalanced medical reports. Our method is a cohesive integration of four modules, which contribute from different perspectives with close collaboration, emphasizing an entity-centered and context-aware design. We confront the linguistic challenge of complex medical reports by distilling precise entity-centered knowledge using a large language model. Subsequently, the catastrophic context-imbalanced issue is addressed by proposing a novel re-balancing factor, then entity-centered descriptors masking is proposed to strengthen the entity-centered context for masked language modeling. Furthermore, a context-guided super-resolution learns to capture the critical pathology vision features with entity-awareness. In addition, multi-scale context fusion helps to optimize informative global and local representations simultaneously. These four readily accessible modules form the core of our simple-yet-effective general Med-VLP framework, referred to as ECAMP, whose significant superior performance is validated on various public datasets across CXRs and fundoscopies on downstream tasks, including classification, segmentation, and detection. ECAMP facilitates the development of a robust medical foundation model, ultimately alleviating the workload of experts. However, the absence of explicit alignment between text and vision modalities hinders the application in zero-shot setting. In the future, we plan to incorporate contrastive learning to further empower ECAMP to function as an expert doctor without any manual annotation.
7 Acknowledgments
This work was supported by National Natural Science Foundation of China under Grant 62271465, Suzhou Basic Research Program under Grant SYG202338, and Open Fund Project of Guangdong Academy of Medical Sciences, China (No. YKY-KF202206).